According to the following reaction, how many grams of ammonia will be formed upon the complete reaction of 24.0 grams of ammonium chloride?

ammonium chloride (aq) = ammonia (g) + hydrochloric acid (aq)

grams ammonia
Answer #1


Molar mass of NH4Cl,
MM = 1*MM(N) + 4*MM(H) + 1*MM(Cl)
= 1*14.01 + 4*1.008 + 1*35.45
= 53.492 g/mol

mass of NH4Cl = 24 g
mol of NH4Cl = (mass)/(molar mass)
= 24/53.49
= 0.4487 mol


Balanced chemical equation is:
1NH4Cl ---> 1NH3 + 1HCl


According to balanced equation
mol of NH3 formed = moles of NH4Cl
= 0.4487 mol


Molar mass of NH3,
MM = 1*MM(N) + 3*MM(H)
= 1*14.01 + 3*1.008
= 17.034 g/mol


mass of NH3 = number of mol * molar mass
= 0.4487*17.03
= 7.643 g
Answer: 7.64 g